5. Feral pigeons (Columba livia domestica) have long been considered the most frustrating pest bird. Flocks of these pigeons# individuals # Z alleles #Palleles #alleles Genotypic Frequency ZP PP Total C) What are the allele frequencies for the two aHow to complete this table ZZ represents p2, PP represents q2 and ZP represents 2pq.
Answer #1

ZZ individuals have highest egg viability. Number 20

ZP individuals have intermediate egg viability.Number=100-(60+20)=20

PP individuals have low egg viability. Number =60

One individual has two alleles.

Please look in the table;

# individuals # Z alleles #Palleles #alleles Genotypic Frequency ZP 20 0.2 140 40. 0 40 20 0.2 20 20 40 60 | 0.6 | 60 160 120

C) The allelic frequentcy for Z allele is 0.3

Genotypic frequency of ZZ + (1/2)× genotypic frequency of ZP= 0.2 +(1/2)×0.2= 0.3

Allelic frequency of P is 1- Z=1-0.3=0.7